OpenStudy (anonymous):

3 people think of a number from {1,2,3,4,5,6,7} show that the probability that a) the 3 integers are different, given that the least integer selected is 5 equates to 6/19 b) the sum of the 3 integers is more than 15, equates to 56/343 ** I have reworded the question as the previous answers were not correct

OpenStudy (dumbcow):

I get different answers: a) 1/7 b) 5/21

OpenStudy (anonymous):

one approach that i have taken is: 1. given that one person has selected 5 then picking the 6 is 1/6 and picking the 7 is 1/7 so that equates to 1/49. 2. however there is no stipulation on the order so the there are 3! ways to pick the 5,6,7 so we get 6/49.

OpenStudy (anonymous):

could the 6/19 be a typo?

OpenStudy (dumbcow):

i dunno, i assumed order didn't matter

OpenStudy (anonymous):

up till now my text has been very accurate, but that question has me stumped

OpenStudy (anonymous):

any idea about part b

OpenStudy (dumbcow):

there are 63 total combinations 7C1 + 7C2 + 7C3 = 7 + 21 + 35 = 63

OpenStudy (anonymous):

for part b) i attempted it like so: all unique numbers summing > 15 , max 7 7+6+5, 7+6+4, etc and progressed through those, each choice can be arranged 3! ways and then went for double entries like 7+7+5 etc that can be arrange 3 ways added then up I think that works but it took a long time there has to be a better way!!!

OpenStudy (dumbcow):

Only 15 of which have sum greater than 15: 666 777 367 467 567 556 557 664 665 667 772 773 774 775 775

OpenStudy (anonymous):

missing 754

OpenStudy (anonymous):

some of them can be chosen in multiple ways making them more likely so 367,467,754,567 can each be selected in 3! or 6 ways giving 24

OpenStudy (dumbcow):

oh yeah, so its 16/63 then ..i thought order didn't mat

OpenStudy (anonymous):

there are 5 with 2 of the same digits which can be selected each in 3 ways giving another 15

OpenStudy (anonymous):

sorry 7 with 2 digits so thats 21

OpenStudy (dumbcow):

hey thats 56/343 i count 10 with 2 digits...thats 30 plus 24 plus 2

OpenStudy (anonymous):

plus the 2 with unique numbers totals 24+21+2 out of 7*7*7 gives us 47/343

OpenStudy (anonymous):

hmmm i think we must be missing some

OpenStudy (anonymous):

765 764 763 754 each avail in 6 combinations = 24 776 775 774 773 772 766 755 655 665 664 each avail in 3 combinations = 30 666 777 each availin 1 combination = 2 out of 7*7*7 = 343 equals 56/343

OpenStudy (anonymous):

However there has to be a much better way than that, it was a past paper question and that took ages!

OpenStudy (anonymous):

Please correct my justification for part a) The total ways of getting 5,6,7 in any order is 6 The total ways of selecting any three numbers (inc. repetitions) is 7^3 So why not 6/343

OpenStudy (anonymous):

the question says that 'given that 5 is the least selected number' - could that be it? I find the terminology is confusing !

OpenStudy (anonymous):

I think that means that 5 is given as selected I also guess that 5 doesn't have to be the first selected?

OpenStudy (anonymous):

so we are really saying what is prop of selecting 6 and a 7 out of the selection, but the 5 could be in any order

OpenStudy (anonymous):

** that is how i read it